OpenStudy (anonymous):

What is the simplest form of the radical expression? Medal and fan for help.

OpenStudy (anonymous):

\[4^{4}\sqrt{2x}+6^{4}\sqrt{2x}\]

OpenStudy (anonymous):

Since\[\sqrt{2x}\] Is in both terms... \[4^{4}+6^{4}(\sqrt{2x})\] Therefore... \[=1552\sqrt{2x}\] Now it has been reduced from two terms to one term. :)